### v1

# A Mapping Analysis of Requirements Between the CRA and the GDPR

Abstract
A new Cyber Resilience Act (CRA) was recently agreed upon in the European Union (EU). The paper examines and elaborates what new requirements the CRA entails by contrasting it with the older General Data Protection Regulation (GDPR). According to the results, there are overlaps in terms confidentiality, integrity, and availability guarantees, data minimization, traceability, data erasure, and security testing. The CRA’s seven new essential requirements originate from obligations to (1) ship products without known exploitable vulnerabilities and (2) with secure defaults, to (3) provide security patches typically for a minimum of five years, to (4) minimize attack surfaces, to (5) develop and enable exploitation mitigation techniques, to (6) establish a software bill of materials (SBOM), and to (7) improve vulnerability coordination, including a mandate to establish a coordinated vulnerability disclosure policy. With these results and an accompanying discussion, the paper contributes to requirements engineering research specialized into legal requirements, demonstrating how new laws may affect existing requirements.
Index Terms:
Legal requirements, essential requirements, cyber security, regulations, compliance, conformance, redundancy
I Introduction
The paper examines the legal requirements imposed by the EU’s CRA and GDPR regulations.111 Regulation (EU) 2024/2847 and Regulation (EU) 2016/679, respectively. The former was agreed upon in December 2024, while the latter came enforceable in 2018. As two regulations are considered, a natural motivation is to ask (Q.1): what is common between them? This motivation aligns well with requirements engineering because overlapping requirements, whether due to a repetition by different stakeholders or something else, is a classical topic in requirements engineering [5, 31]. Among other things, overlapping requirements may cause redundancy; therefore, the motivation can be sharpened by considering an imaginary software or system that is already compliant with the GDPR’s legal requirements. Then, after it is known what is common, it can be deduced what is new. In other words (Q.2): what new legal cyber security requirements the CRA brings?
Redundancy is a classical concept also in legal scholarship and related fields [13, 21]. The reason is understandable: laws often repeat themselves. This legal repetition is partially explained by lawmaking, which is often an iterative endeavour in a sense that new laws often (but not always) build upon existing laws. The concept of an architecture provides a related analogy; the “high-dimensionality of the architecture of both law and computer science implies that choices made at any point of the system will ripple through the entire system, resulting in bugs or new features” [14, p. 2]. Given this insightful quotation, the focus is on new requirements but not on new features; the scope is restricted to requirements engineering alone, without delving into a question on how these might be designed and implemented, or, given the earlier remark about some existing GDPR-compliant implementation, how it should be changed. Nevertheless, the quotation’s point about ripple effects provides another motivation for the paper. As was envisioned in an early requirements engineering research regarding the GDPR [15], and as has also generally been observed  [26, 33], also legal requirements change over time, whether due to changing interpretations and guidelines by regulators, court practice and case law, enactment of new laws or deprecation of existing laws, or something else.
Figure 1: Changes, Ripples, and the Paper’s Scope
Overlaps, redundancy, duplication, and related terms justify also what is meant by a mapping analysis in the present context; the focus is on the intersection of the legal requirements imposed by the two EU regulations on one hand, and the relative complement of the requirements on the other hand. The latter refers to the CRA’s new requirements that are not also a part of the GDPR. By using this set theory terminology, a necessary framing follows: (a) the focus is on implementations that process personal data (see Fig. 1). In this regard, it should be emphasized that the CRA applies also to non-personal data. This framing may also explain the legal redundancy; strict references to the GDPR are not possible in most cases because the CRA’s scope is much wider in terms of cyber security. A further framing is important to mention: (b) the CRA is only considered in terms of its so-called essential cyber security requirements—a term elaborated later in Subsection III-C. Furthermore, (c) the GDPR is only considered in terms of its intersection with the CRA. In light of the existing and related work briefly outlined in the opening Section II, this choice is justifiable due to the vast amount of research on the GDPR. There is no reason to commit to redundancy also in science.
II Related Work
The GDPR has received a considerable attention after its enactment, including in requirements engineering and associated computing disciplines [2, 15, 30, 44]. Due to the CRA and the mapping approach, the related work also expands toward cyber security research. In this regard, it can be mentioned that cyber security violations rank high also with respect to GDPR enforcement fines, although still lower than the regulation’s basic legal principles [40]. Such cyber security violations likely also continue today and in the nearby future because (personal) data breaches and ransomware operations are currently the two most pressing high-impact risks for many organizations [9]. These also illustrate the basic concepts discussed later in Subsection IV-A: at minimum, the former typically violate confidentiality and the latter availability.
Due to other recent ripples in the EU’s legal framework for cyber security, the GDPR has also been considered in conjunction with other recent cyber security laws [28]. Thus far, however, existing research on the CRA is scant but not entirely absent, the likely explanation being simple; at the time of writing, the regulation is new and the details are still being worked on in the member states. Regarding research, besides legal overviews [29], the CRA’s new requirements for vulnerability coordination and disclosure have been recently addressed [43]. There is also work on the CRA’s still unclear implications upon open source software [4]. The regulation’s requirement about secure default configurations has also been implicitly addressed through a literature review [38]. These and potentially other outlying examples notwithstanding, the volume of existing research is still low. For this reason, the CRA also takes a priority in the mapping in the form of Q.2. Further related work can be pointed out by considering some of the core concepts together with additional framings.
III Concepts and Further Framings
III-A Cyber Security and Data Protection
Cyber security and data protection have an intrinsic relationship already because confidentiality belongs to the former. Indeed, the GDPR contains also explicit obligations to protect personal data of natural persons from cyber security threats. When considering the intersection of these two regulations, the framing sharpens further toward the GDPR’s cyber security obligations; the regulation’s many other requirements are framed out. While cyber security is fully in the CRA’s scope, the framing toward existing but imaginary implementations further restricts the mapping toward technical aspects, which are also what the essential requirements are about. Because the CRA is a product-specific regulation, whereas the GDPR applies to anything processing personal data, including organizations, the framing is toward products that process personal data. Such products must comply with both regulations.
As for cyber security, it can be noted that the concept is much broader than older related concepts, such as information security within which information (including personal data) is the asset to be protected. In both regulations, but particularly in the CRA, the broader scope is seen in the use of the resilience concept, among other things. As will be seen, there are also new mandates that consider the availability of other networks and services as an asset that should not be harmed by products.
III-B Requirement Types
There is a classical distinction between functional and non-functional requirements in software engineering. The former refer generally to something that a software should do, whereas non-functional requirements are more general quality attributes of the software, which also constraint the software [18]. Both security and privacy have usually been categorized as non-functional requirements [16, 20, 24]. The same applies to data protection. While some have seen the GDPR as a series of non-functional requirements [22], others have perceived it to contain both functional and non-functional requirements [15, 30]. Similar ambiguities apply in terms of cyber security. For instance, once broken down to smaller details, encryption can be seen to contain both functional and non-functional requirements [12]. By following this line of reasoning, the mapping analysis approaches the distinction from a perspective of composition; the overall non-functional requirement of cyber security can be seen to contain various functional requirements describing what an implementation should do. These functional requirements are understood as being representable in a flow diagram operating at the same abstraction level as a given non-functional requirement, which, in contrast, is not a process. With respect to functional requirements, a further separation is done between functional requirements for software, systems, and other implementations and “organizational” requirements with clear functions. As soon discussed in Subsection III-D, the distinction is relevant because the CRA’s functional requirements are dependent on the organizational (functional) requirement of risk analysis.
A further clarification is required: the paper operates with legal requirements, which are generally about something that a law requires for software, systems, and other implementations. The term legal requirement is important for three reasons. First, it separates law-imposed requirements from “conventional” requirements, such as those elicited from clients and customers. With respect to non-functional requirements in general, legal requirements are particularly something that are non-negotiable [47]. Second, compliance with legal requirements can be seen as an overall non-functional requirement from which, if possible, functional requirements should be drawn. As non-functional requirements too should be verifiable [47], this point reiterates the point about processes. Third, legal requirements are also a distinct genre within requirements engineering research. Despite various models, frameworks, techniques, and tools, the challenges in this genre are well-recognized [15, 33]. In what follows, no specific methodology is required because the GDPR’s requirements have already been addressed in previous research, and the CRA’s requirements are relatively straightforward to identify (but not necessarily interpret) because the noted framing toward the regulation’s essential requirements. This term needs a clarification too. It also serves to pinpoint a few brief points about compliance and conformance with the CRA regulation.
III-C Conformance and Compliance
The CRA categorized products into three groups: “normal” products (Article 6), important products (Article 7), and critical products (Article 8). The “normal” ones must comply with the essential cyber security regulations enumerated in the regulation’s Annex I. These are also the paper’s focus. Then, important products, which are listed in Annex III, must further comply with conformity assessment procedures specified in Article 32. While there are different options in these procedures, a typical choice for many vendors will likely be a self-assessment of conformity, as has also been common with the EU’s product safety laws [37]. It is also possible to choose a stricter procedure involving third-party audits. Such audits are required from the critical products. These are linked to the cyber security certification scheme established in the Cybersecurity Act, that is, Regulation (EU) 2019/881. The audits themselves will be supported by standards [11, 29]. For the paper’s purposes, it is important to emphasize that, with a few exceptions, practically all information technology products need to comply with the CRA’s essential requirements.
III-D Requirements and Risks
Both the CRA and the GDPR are risk-based regulations, meaning that measures taken, including requirements elicited from the regulations, should be balanced against results from risk analyses. Given the exhaustive body of knowledge on cyber security risk management [46], which aligns with the CRA’s obligations, it suffices to remark that the GDPR’s risk analysis mandates, which include privacy impact assessments, can be done analogously by considering risks to natural persons specifically. In fact, classical risk matrices involving impacts and probabilities have been recommended in the GDPR context [7]. This point is important to emphasize because risks are a classical way to prioritize security requirements, including countermeasures and security design solutions in general [27]. Although such a prioritization is left for further work as it requires an existing implementation or at least a design for it, it should be understood that not all of the legal requirements carry the same weight in each and every case.
IV Overlaps
In what follows, the overlapping legal requirements are discussed and elaborated. The overlaps should be understood to apply with respect to the framings noted in the introduction. Before continuing, therefore, it should be understood that the CRA applies to all product with a “a direct or indirect logical or physical data connection to a device or network”, as stated in Article 2(1). Thus, anything with a network connection is in the scope. The notable exclusions are maritime equipment and ships, medical devices, motor vehicles, and cloud computing, including software, platform, and infrastructure as a service. The likely reason is that these excluded domains are covered in their own specific laws. With the framing toward the GDPR, a context of the overlaps could thus range from Internet of things (IoT) devices to information systems and beyond.
IV-A The CIA Triad
The classical confidentiality, integrity, and availability (CIA) triad appears in both regulations (see Fig. 2). Like cyber security itself [20], the CIA triad has usually been interpreted as a non-functional requirement [24]. This interpretation seems suitable also for the present work because neither of the two regulations define any specific technical functionality on how to guarantee confidentiality, integrity, and availability, the exception of encryption perhaps notwithstanding.
Figure 2: Overlaps Between the CRA’s and the GDPR’s Requirements
With respect to the GDPR, the CIA triad appears literally in paragraphs (b) and (c) in Article 32. In addition to mentioning the triad explicitly, the regulation also obliges to ensure resilience, which includes an ability to restore availability and access to personal data during incident management. The article’s paragraph (a) further mentions encryption and pseudonymization, which are both related to confidentiality. As the latter concept is specific to the GDPR, it is omitted from the mapping analysis without a notable loss of specificity.
Then, regarding the CRA, the CIA triad is addressed through four essential requirements. Of these, the confidentiality and integrity requirements use a wording about “stored, transmitted or otherwise processed”. With respect to data, whether personal or non-personal, both of which are in the CRA’s (but not in the GDPR’s) scope, the wording reflects the commonplace security taxonomy of “data at rest”, “data in transmission”, and “data in processing” [6, 19]. Encryption is a typical security requirement for the first two concepts.
With respect to confidentiality, encryption is again explicitly mentioned in the CRA analogously to the GDPR. However, regarding integrity, the scope is wider because the integrity guarantees should cover also “commands, programs and configuration” against manipulation or modification by unauthorized parties. Of these three concepts, unauthorized commands could be reflected against structured query language (SQL) injections or even buffer overflows, while unauthorized programs might cover an execution of unauthorized third-party code, as has been common in the web domain [42], among other domains. Furthermore, there is a specific essential requirement about ensuring a protection against unauthorized access, which too is related to confidentiality and integrity. Authentication and identity management are mentioned as examples.
Availability too should be guaranteed, as always, and again resilience is mentioned also in the CRA as a concept to address post-incident recovery solutions together with mitigative measures against denial-of-service (DoS) attacks. In addition, there is an interesting and relevant new requirement, as specified in paragraph 2(i) in Part I of Annex I, to minimize any negative effects upon products and connected devices themselves “on the availability of services provided by other devices or networks”. This essential requirement can be reflected against a concept of cascading risks, failures, and incidents; a single event may lead to further events, whether in terms of cyber security or more generally [8, 23]. A prime historical example would be the Mirai botnet that was built upon compromised IoT devices [25]. It remains to be seen whether the CRA may decrease a probability of such botnets emerging in the future, and whether a producer or producers could be held liable for enabling such emergencies through gross negligence of the essential cyber security requirements.
IV-B Data Minimization
Data minimization is an important principle and an obligation in the GDPR; one should only collect a minimal amount of data necessary to fulfill a given purpose. Interestingly enough, the CRA specifies the same obligation; only what is necessary to a purpose should be collected. The mandate is interesting because the CRA applies also to non-personal data. In any case, data minimization should be balanced with traceability.
IV-C Traceability
Both regulations impose a traceability requirement. Although logging is the obvious solution to fulfill the requirement [15], the term traceability in itself is slightly better because the GDPR applies also to manual processing of personal data. In any case, the CRA’s paragraph 2(l) in Part I of Annex I requires “recording and monitoring relevant internal activity, including the access to or modification of data, services or functions”. Thus, not only should the GDPR-specific operations, such as an erasure or a rectification of personal data, be logged but the logging should in principle apply to all CRUD operations, to use the common abbreviation for create, read, update, and delete. As said, however, balancing should be likely still done in many cases, and also for scalability reasons [36]. A further point is that the wording used about an access to services and functions indicates also other considerations. To use the IoT domain again as an example: a device’s connection to a network might be logged.
Figure 3: The CRA’s New Essential Requirements
IV-D Data Erasure
Although the right to personal data protection is not absolute, the GDPR grants a right to delete one’s personal data in many cases. Also the CRA contains this requirement but with a more general wording about a secure and easy removal “on a permanent basis all data and settings”. Although settings should be deleted too, the term data erasure is used in Fig. 2 due to the framing. The emphasis on permanent and secure deletion reiterates a classical topic in cyber security [34]. The erasure requirement is also generally related to the CRA’s mandate to support products for a minimum of five years in most cases. The relation between life cycles and the erasure requirement comes from a broader contemplation about sustainability, the green transition, and second-hand markets.
IV-E Security Testing
Figure 4: Non-Functional, Functional, and Organizational Requirements Identified from the CRA’s Essential Cyber Security Requirements
Both the CRA and the GDPR require security testing. In terms of the latter, the requirement is specified in Article 32(d), which states that there should be “a process for regularly testing, assessing and evaluating the effectiveness of technical and organisational measures for ensuring the security”. The coverage of organizational security reflects the conceptual issues that were noted in Subsection III-A. In the CRA a wording “effective and regular tests and reviews” is used. In other words, also security reviews should be done alongside testing.
V The CRA’s New Essential Requirements
The new cyber security requirements imposed by the CRA regulation can be grouped into seven groups. These are summarized in Fig. 3 and briefly elaborated in what follows. To summarize the points raised, Fig. 4 additionally displays the functional requirements explicitly identifiable from the CRA. While risk analysis is omitted from the figure for brevity, it too is a functional and organizational requirement. In addition, many of the CRA’s essential requirements, including the earlier ones in Fig. 2, must be documented according to Article 13.
V-A No Known Vulnerabilities
The first new essential requirement is simple enough: all products must “be made available on the market without known exploitable vulnerabilities”. The wording about exploitability gives some leeway for vendors, but otherwise the legal mandate is on the side of straightforward functional requirements. To be able to fulfill the requirement, a vendor obviously must know which components it uses, whether software or hardware components. In terms of software, the requirement therefore aligns with a mandate to establish a software bill of materials for each product. There is also a philosophical angle to the requirement: known by whom? While the regulation’s future guidelines and its enforcement may clarify the question, presently it would seem sufficient to focus on the major international vulnerability databases.
V-B Secure Defaults
The CRA mandates that all new products entering the EU’s internal market should be shipped with secure default configurations. Though, an opt-option is given with a wording “unless otherwise agreed between manufacturer and business user in relation to a tailor-made product”. In any case, as this requirement has been discussed in detail recently [38], it suffices to remark that the requirement is a classical security design principle. For instance, the OpenBSD operating system has long endorsed the principle [32]. Having said that, the IoT domain’s track record exemplifies that the principle has often been also ignored or even abused [3]. Therefore, the CRA’s new essential requirement can be generally welcomed.
V-C Security Updates
All products should be generally supported for a minimum of five years according to the CRA. Though, the wordings in recital 60 and Article 13(8) are vague, suggesting that a shorter support period is also possible in some cases. Regardless, the essential requirement is that security patches for vulnerabilities should be provided free of charge throughout a product’s life cycle. Furthermore, vulnerabilities should be addressed “where applicable, through automatic security updates”. Even when applicable, the automated updating functionality should come with “a clear and easy-to-use opt-out mechanism” as well as notifications to users about updates and a further option to “temporarily postpone them”. Regardless whether automation is used or not, updates should be securely distributed. Furthermore, “where technically feasible, new security updates shall be provided separately from functionality updates”. These clarifications and options are summarized in Fig. 5.
Even though optional, automatic updates can be seen as particularly relevant for improving the IoT domain’s cyber security posture [3]. The provisions and options provided also indicate that the CRA seems to have taken into account most—or even all—of the deadly sins associated with software updates, as elaborated in the literature [17]. However, the CRA’s wide scope, which includes everything from smartphones to industrial control systems, will likely bring new challenges for some vendors. In addition, there are always risks involved too in updating; in fact, annual surveys indicate that faulty updates have been relatively common in some industry sectors [10]. Regressions are also unfortunately common in many software projects, including critical ones, such as the Linux kernel [39]. These and other potential problems warrant also further research in software engineering and beyond.
Figure 5: Updating Requirements
V-D Attack Surface Minimization
There is also an essential and explicit requirement to minimize attack surfaces. The CRA mentions external interfaces in this regard, but nothing beyond that is available. Although there is plenty of research in this area, including in the IoT context [45], a systematization of knowledge might be beneficial now that the topic is covered also in a new law. It can be also noted that this requirement seems to contradict slightly the requirement of security updates. In other words, the functionality in Fig. 5 also enlarges attack surfaces.
V-E Exploitation Mitigation
The CRA requires that appropriate exploitation mitigation techniques should be designed, developed, and enabled. This essential requirement aligns partially with a security design principle of turning security features on [38]. However, the CRA does not say anything specific about the techniques.
V-F SBOMs
As already noted, SBOMs must be implemented for all products. While useful also for vendors and their supply-chain management, they may be used for enforcement purposes too, given that a vendor is obliged to deliver its SBOMs to supervising authorities upon request according to Article 13(25). The Article’s context and motivation are about potential large-scale sweeps of open source software dependencies. Confidentiality should be guaranteed when dealing with authorities and vendors are not obliged to make their SBOMs public.
V-G Vulnerability Coordination
Finally, the CRA brings some new requirements for vulnerability coordination, including vulnerability disclosure. As this requirement has already been considered in existing work [43], it suffices to merely enumerate some of the particular requirements: vulnerabilities should be fixed without delay, updates fixing them should be disseminated without delay, they should be documented in conjunction with SBOMs, security advisories should be delivered to users alongside public disclosures, information sharing about vulnerabilities should be facilitated, coordinated vulnerability disclosure policies should be established and followed, and so forth and so on.
VI Discussion
VI-A Conclusion
The paper examined the new CRA regulation by comparing and contrasting it with the older GDPR. While the CRA’s scope is much wider, including most products with a networking functionality and covering also processing of non-personal data, the mapping study conducted revealed also similarities. Notably, the CIA triad guarantees, data minimization, traceability, data erasure, and security testing appear in both regulations, although a verbatim equivalence is not present (Q.1). Regarding non-overlapping requirements, the CRA imposes mandates to launch products without known vulnerabilities and with secure default configurations, to provide security updates, to minimize attack surfaces, to enable appropriate exploitation mitigation techniques, to implement SBOMs, and to follow established vulnerability coordination practices (Q.2). All in all, the requirements can be seen as sensible for the regulation’s goal of improving cyber security in Europe. Nor do these seem overly burdensome to design and implement, although the CRA’s broad scope may cause some problems in some industry sectors with specialized products.
VI-B Further Overlaps
As has been pointed out [29], there is a further overlap between the two regulations and the radio equipment Directive 2014/53/EU. There are thee obligations specified in the directive’s Article 3(3)(d), (e), and (f) that overlap with the CRA and the GDPR: personal data should be protected, fraud prevention should be implemented when possible, and an “equipment does not harm the network or its functioning nor misuse network resources, thereby causing an unacceptable degradation of service”. The obligation quoted resembles the point about cascading effects noted in Subsection IV-A.
As has again been pointed out [41], the GDPR also overlaps slightly with the EU’s recent Directive (EU) 2022/2555, which is commonly known as the NIS2 directive. The overlap comes from incident notifications about serious personal data breaches to public authorities. Although a direct overlap is not present, also the CRA imposes incident notification obligations. The regulation’s Article 14 mandates manufacturers to report about actively exploited vulnerabilities as well as severe incidents to the supervising authorities. While the deadlines of reporting align with those specified in the NIS2 directive, the definitions for severe incidents are different. According to the CRA’s Article 14(5), these refer to those that affect the CIA guarantees of products on one hand, and those “leading to the introduction or execution of malicious code in a product with digital elements or in the network and information systems of a user of the product” on the other hand. Once again, the theme of cascading effects is present in this direct quotation.
VI-C Further Work
With the results from the mapping analysis, the paper contributes to the early efforts to build a reusable body of knowledge on legal requirements [33]. Further work is required to go beyond requirements; to evaluate what must be done or changed in the domain of actual software, systems, and other implementations. However, due to the CRA’s broad scope, one-size-fits-all solutions or recommendations are presumably difficult to postulate. Thus, case studies would offer a good path forward to better contextualize the requirements elicited from the regulation. Another option would be to maintain a high level of abstraction and evaluate how the CRA’s requirements align with existing cyber security design principles. For instance, principles related to secure defaults, fails-safes, access controls, and the zero trust model [38] seem to align quite well with the CRA’s new requirements.
Further work is required also on risk analysis and auditing. While the details are still being worked on in the EU, a good topic related to risk analysis would be an evaluation of the suitability of existing standards for auditing, whether internally or externally, a fulfillment of the CRA’s essential requirements. Although there is existing work in this area [11], risk analysis itself seems to have received less attention. As was noted in Subsection III-D, risk analysis should guide the prioritization of the essential requirements and their implementations.
Auditing would be a good topic also otherwise because cyber security audits are only rarely conducted according to some industry surveys [35]. Regarding the survey results referenced, it seems that the CRA patches also some other notable limitations in industry practices. Although costs are certainly a factor, also attack surface analysis and security reviews more generally seem to be only occasionally done in the industry, among other things. In light of such observations, another industry survey about the CRA specifically would also shed light on the potential challenges ahead regarding compliance. As already remarked, it be that the challenges are sectoral, specific to some particular products, or both. For instance, industrial control systems and supervisory control and data acquisition (SCADA) systems have been seen to suffer from many of the issues addressed by the CRA. That is, among other things, these systems do not always have robust authentication and authorization methods, remote access contains particular risks, updates are often lacking, and monitoring and logging are oftentimes insufficient [1]. These and other real or perceived challenges correlate with real or perceived problems in addressing them. For instance: due to the criticality of the systems in many industry sectors, including those related to countries’ critical infrastructures, integrity guarantees and pre-delivery and post-delivery verification of updates are likely of significant importance. That is, faulty updates alone are likely a high-impact risk with these systems and in these sectors.